Cos 30 degrees = Sin ??? degrees
Answer by mangopeeler07(462)

cos (x)=sin(90-x)
cos (30)=sin (90-30)
cos (30)=sin (60)

Why?
Because cos is adjacent over hypotenuse
and sin is opposite over hypotenuse

If you draw a 30, 60, 90 degree triangle, the adjacent for 30 becomes the opposite for 60. So the cos of 30 would be the sin of 60. So it is the same ratio by a different name.